OpenStudy (anonymous):

the expression in radical notation 81^-3/4

OpenStudy (anonymous):

81^(-3/4) = 1/ 81^(3/4) = 1/[81^(1/4)]^3 = 1/(3^3) = 1/27

OpenStudy (anonymous):

radical notation is\[\sqrt[4]{81^{-3}}\]

OpenStudy (anonymous):

so is the evaluated? I need to know how to write it in radical notation first

OpenStudy (anonymous):

See my last post.

OpenStudy (anonymous):

ok so then when it is evaluated the answer would be 1/27

OpenStudy (anonymous):

yes. That is the fractional answer as evaluated. The radical notation is hard to work with so it is not a "final answer".

OpenStudy (anonymous):

Making sense?

OpenStudy (anonymous):

thanks!!